# Speech Robust Bench: A Robustness Benchmark For Speech Recognition

### v2

@@ 3.3 Metrics
SRB measures the utility of the model with the widely used Word Error Rate. WER is computed as the word-level edit distance between the reference and the predicted transcripts, normalized by the length of the reference (see Appendix B for formal definitions). To measure the robustness of the model to input perturbations, we use WER Degradation (WERD), computed as $WER(\mathcal{X}_{p})-WER(\mathcal{X})$, where $\mathcal{X}$ contains clean audio, while the audio in $\mathcal{X}_{p}$ is perturbed by perturbation $p$. If $p$ is synthetically applied, both $\mathcal{X}$ and $\mathcal{X}_{p}$ contain the same utterances, otherwise $\mathcal{X}$ contains clean audios from LibriSpeech and TEDLIUM.
When aggregating metrics (WER/WERD) over multiple perturbations, we follow the practice of [1] and divide the metric by a measure of difficulty, i.e., by the (estimated) speech quality degradation. This adds weight to errors on “easy” corruptions (less quality degradation) and under weighs errors on “harder” corruptions (more quality degradation) when computing averages. We refer to the difficulty normalized versions of WER/WERD as Normalized WER/WERD (NWER/NWERD). We estimate speech quality using DNSMOS [60] and PESQ[61], which are models of human judgments of speech quality and predict Mean Opinion Scores (MOS) [62]. PESQ uses various signal processing methods to predict MOS, while DNSMOS uses DNNs to do the same. To compute speech quality degradation we compute PESQ and DNSMOS for each clean and noisy recording multiplied by -1 (lower values indicate less degradation). Since we are only interested in the relative degradation between perturbations, we normalize the scores to have mean 50 and standard deviation 25.
@@ 4 Evaluation
We evaluate several recent and popular ASR DNNs (§4.1) using SRB and analyze the results to uncover fine-grained differences in their robustness to non-adversarial and adversarial input perturbations. We further extend our analysis by measuring ASR model robustness for various sub-groups, namely English speech and non-English (Spanish) speech, and male and female speakers. Prior works [63, 64] observe that there is a disparity in transcription accuracy between subgroups. Our analysis augments these observations by showing that inter-group disparities in robustness may also exist, thus demonstrating the utility of SRB in the broader field of trustworthy AI.
@@ 4.1 Models
For English speech, we evaluate Whisper[19] variants: large-v2 (wsp-lg), medium (wsp-md), small (wsp-sm), base (wsp-bs), and tiny (wsp-tn, wsp-tn.en), Wav2Vec-2.0 [38] variants: base (w2v2-bs), large (w2v2-lg), self-trained large [65] (w2v2-lg-slf), and Robust Wav2Vec [22], HuBERT [39] variants: large (hubt-lg) and XL (hubt-xl), Nvidia Canary [25] (cnry-1b), MMS [28] (mms-1b), Speech-T5 [66] (spch-t5) and DeepSpeech [67] (ds). For Spanish speech, we evaluate on the mono-lingual Wav2Vec base Spanish [68] (w2v2-bs-es), Wav2Vec XLSR Spanish [69] (w2v2-lg-es), wsp-lg, wsp-bs, wsp-tn, and mms-1b. We used the Huggingface implementations for all the models except cnry-1b (https://huggingface.co/nvidia/canary-1b) and ds (https://github.com/SeanNaren/deepspeech.pytorch). More details about the models are in Table 3 in the Appendix.
4.2 Accuracy on Clean Data
Figures 3(a) and 3(d) show the WER of the models on clean English and Spanish, respectively. We see that cnry-1b is the most accurate model on both languages, followed by w2v2-lg-slf on English and wsp-lg on Spanish.
4.3 Robustness of English ASR Models to Non-Adversarial Perturbations
Fig. 3(b) presents the NWERD for the English ASR models under non-adversarial noises and perturbations. We observe that wsp-lg emerges as the most robust model, followed by wsp-md and cnry-1b. Despite having the lowest overall WERD, wsp-lg is outperformed by wsp-md and cnry-1b on white noise, environmental noise, and special effects (see Appendix D for more detailed results). The massively multilingual mms-1b has the highest NWERD in social scenarios and far-field recordings category, indicating that it is unsuited for applications like voice assistants and smart home set-ups. We also notice that Wav2Vec-2.0 based models also struggle with the “speech variations” indicating that they are not suitable when users have diverse accents. Surprisingly, spch-t5 is the least robust model, and performs even worse than the much smaller and older ds. Similarly, wsp-md outperforms the newer and larger cnry-1b on almost all corruption categories.
@@ 4.3 Robustness of English ASR Models to Non-Adversarial Perturbations
Figure 4: WERD of cnry-1b, wsp-md and wsp-sm on perturbations in the spatial acoustics and special effects categories.
@@
Upon closer examination, we see that cnry-1b has higher NWERD than its peers (wsp-md and wsp-sm) for two categories of perturbations: spatial acoustics and special effects. Upon looking at the WERD on each perturbation within these categories (Fig. 4) we find that cnry-1b is more sensitive than its peers to echo, real room impulse responses, and speed and pitch modifications. This analysis demonstrates that SRB can evaluate the robustness of ASR models at multiple ganularities and can pinpoint the weaknesses of a given model. Detailed results can be found in Figs. 7 and 9 in the appendix.
@@
Takeaway: cnry-1b is the most accurate, but not most robust; Wav2Vec variants struggle against speech variations and social settings, thus, may not be suitable when users have diverse accents.
4.4 Robustness of English ASR Models to Adversarial Perturbations
@@
Fig. 3(c) shows the WERD of the models on utterance specific and utterance agnostic adversarial attacks. We observe that hubt-lg and hubt-xl emerge as the most robust models to utterance-specific adversarial attacks, followed by other models based on Wav2Vec-2.0 large, Whisper, and Canary. Interestingly, Wav2Vec models that exhibited mediocre robustness to non-adversarial perturbations are robust to utterance-specific attacks, while Whisper models and Canary, which were the most robust on non-adversarial perturbations, struggle on utterance specific attacks. We also note that most Wav2Vec models are considerably more robust to attacks against TEDLIUM than against LibriSpeech, and the opposite is true for whisper and Canary models. Under utterance-agnostic attacks, the best-performing models are mms-1b, wsp-bs, and wsp-sm. It is interesting to note that the smaller variants of Whisper limit the generalizability across utterances of the adversarial perturbations to a greater extent than their larger counterparts.
@@
Takeaway: Models that are most robust to non-adversarial perturbations, are mediocre against adversarial perturbations; cnry-1b is highly vulnerable to utterance specific attacks.
4.5 Robustness vs. Model Attributes
To determine if the prevailing practice of training DNNs with more parameters on larger datasets is yielding improvements in robustness, we use robust linear regression to fit a line to WERD vs. number of model parameters/size of the training data for the candidate models in Figs. 5(a) and 5(b), respectively. Increasing model size is correlated with improved robustness (lower WERD).To further isolate the impact of the model size we plot the NWERD of models from the same family in Fig. 5(c), which have similar architectures and training datasets. We note that larger models are more robust in the Whisper and Wav2Vec-2.0 families, but, surprisingly, not in the HuBert family. In contrast, increasing training data appears to have only a minor influence on robustness (Fig. 5(b)).
Takeaway: Larger models tend to be more robust, while smaller models, even if they are trained on large datasets, are less robust. This runs somewhat counter to the prevailing wisdom [19, 22].
Figure 5: (a & b) WERD for all models with robust regression fitted line on non-adversarial (blue) and adversarial (orange) perturbations, plotted against (a) number of parameters, and (b) hours of training data. (c) NWERD lineplot with non-adversarial and adversarial perturbations, three families of models (color indicated). Models per family share similar architectures and training datasets. (d) Robustness of multi-lingual English (solid) and Spanish (hashed) models.
4.6 Robustness for Population Sub-Groups
In the preceding analysis, we considered robustness aggregated over the entire population (i.e., dataset). However, populations are generally not homogeneous, and, thus, the robustness of the model may differ on various population sub-groups. Prior works have commonly analyzed sub-group fairness of ASR models by comparing the overall WER for each sub-group on a benchmark dataset [70]. It is possible that models that are fair on average, may not be fair under certain conditions. In the following, we use SRB to uncover and analyze the disparities in the models’ robustness across four sub-groups: English and Spanish speech, and male and female speakers. We find that disparities indeed exist, with multi-lingual models generally being more robust for English than Spanish (Fig. 5(d)), and most models being less robust for females than males.
4.6.1 Robustness on Non-English (Spanish) Speech
The clean accuracy of the Spanish ASR models is presented in Fig. 3(d), and the NWERD under non-adversarial perturbations for each model is shown in Fig. 3(e). We observe, that the two most accurate models, cnry-1b and wsp-lg, are also the two most robust models, but mms-1b, which had fairly high WER, is the third most robust model in terms of NWER. We also notice a general trend that all the models, except cnry-1b, struggle against “spatial acoustics” category, which is rather worrying since the influence of spatial acoustics is ubiquitous in the real-world.
@@ 4.6.1 Robustness on Non-English (Spanish) Speech
Considering adversarial attacks (Fig. 3(f)), we observe that Wav2Vec models exhibit better robustness than Whisper and Canary under both utterance-agnostic and utterance-specific perturbations. This is similar to the trends observed in English speech (Fig. 3(c)), and indicates that perhaps learned front-ends may be more adversarial robust than the spectrograms used by Whisper and Canary. Detailed results can be found in Fig. 8 in the appendix.
@@
We also compare the robustness exhibited by multi-lingual models, wsp-lg, wsp-bs, cnry-1b and mms-1b on English and Spanish. The WERD of these models on both languages is presented in Fig. 5(d). We observe that Whisper models achieve lower WERD on English speech than on Spanish on almost all perturbation categories, while cnry-1b and mms-1b achieve similar WERD on some categories. We also note that the difference in WERD on some common perturbation categories, like environmental noise, and spatial acoustics, is much greater for wsp-lg than for cnry-1b.
@@
4.6.2 Robustness and Gender
To measure the disparity in prediction accuracy across genders (males/females), we compute the log of the ratio of the WERs of the ASR model on female and male speakers. We call this measure the Log WER Ratio (LWERR). A positive value of LWERR indicates that the model is biased against females and a negative value indicates that the model is biased against males.

### v3

@@ 3.2 Metrics
SRB measures the utility of the model with the widely used Word Error Rate. WER is computed as the word-level edit distance between the reference and the predicted transcripts, normalized by the length of the reference (see Appendix B for formal definitions). To measure the robustness of the model under challenging scenarios, we use WER Degradation (WERD), computed as $WER(\mathcal{X}_{s})-WER(\mathcal{X})$, where $\mathcal{X}$ and $\mathcal{X}_{s}$ are datasets containing clean speech and speech from scenario $s$, respectively. For scenarios (1)-(3.1) (see § 3.1), $\mathcal{X}_{s}$ is an inherently noisy dataset, and $\mathcal{X}$ will be LibriSpeech for English and Multi-Lingual LibriSpeech for Spanish. For scenarios (3.2)-(6), $\mathcal{X}$ is a clean dataset, and $\mathcal{X}_{s}$ is a perturbed version of $\mathcal{X}$.
When aggregating metrics (WER/WERD) over multiple scenarios, we follow the practice of Hendrycks & Dietterich (2019) and divide the metric by a measure of difficulty, i.e., by the (estimated) speech quality degradation. This adds weight to errors on “easy” scenarios (less quality degradation) and underweights errors on “harder” scenarios (more quality degradation) when computing averages. We refer to the difficulty normalized versions of WER/WERD as Normalized WER/WERD (NWER/NWERD). We estimate speech quality using DNSMOS (Reddy et al., 2019) and PESQ(Rix et al., 2001; Miao Wang & ananda seelan, 2022), which are models of human judgments of speech quality and predict Mean Opinion Scores (MOS, Rec 2018). PESQ uses various signal processing methods to predict MOS, while DNSMOS uses DNNs to do the same. To compute speech quality degradation we compute PESQ and DNSMOS for each clean and noisy recording multiplied by -1 (lower values indicate less degradation). Since we are only interested in the relative degradation between scenarios, we normalize the scores to have mean 50 and standard deviation 25.
Note on usage: We use NWERD for non-adversarial scenarios (1-5) but WERD for adversarial attacks because adversarial attacks are model-specific and thus DNSMOS/PESQ scores for adversarially perturbed audio will be different for each model, which will lead to a different normalization during NWERD computation and make comparisons difficult.
@@ 4 Evaluation
We evaluate several recent ASR DNNs (§4.1) using SRB and analyze the results to uncover fine-grained differences in their robustness in various challenging scenarios. We further extend our analysis by measuring ASR model robustness for various sub-groups, namely English speech and non-English (Spanish) speech, and male and female speakers. Prior works (Liu et al., 2022; Veliche & Fung, 2023) observe that there is a disparity in transcription quality between subgroups. Our analysis augments these observations by showing that inter-group disparities in robustness may also exist, thus demonstrating the utility of SRB in the broader field of trustworthy AI.
@@ 4.1 Models
For English, we evaluate Whisper (Radford et al., 2023) large-v2, base, medium, small, and tiny (wsp-{lg,bs,md,sm,tn}), Wav2Vec-2.0 (Baevski et al., 2020) base, large, self-trained large (Xu et al., 2021), and Robust Wav2Vec (Likhomanenko et al., 2020) (w2v2-{bs,lg,lg-slf,lg-rob}), HuBERT (Hsu et al., 2021a) large and XL (hubt-{lg,xl}), Nvidia Canary (NVIDIA, ) (cnry-1b), Nvidia Parakeet RNN-T and CTC (NVIDIA, ) with 0.6B and 1.1B parameters (prkt-rnnt-{0.6,1.1}b, prkt-ctc-{0.6,1.1}b), MMS (Pratap et al., 2020) (mms-1b), Speech-T5 (Ao et al., 2022) (spch-t5), DeepSpeech (Amodei et al., 2016) (ds), and Speechbrain (Ravanelli et al., 2024) models with Conformer encoders, and transformer and RNN-T decoders. For Spanish speech, we evaluate mono-lingual Wav2Vec base Spanish (Wang et al., 2021b) (w2v2-bs-es), Wav2Vec XLSR Spanish (Conneau et al., 2020) (w2v2-lg-es), wsp-{lg,bs,tn}, and mms-1b. We used the Huggingface implementations where available, except ds (https://github.com/SeanNaren/deepspeech.pytorch). More details about the models are in Table 5.
clean accent audio proc noise (env) noise (white) sFX social (FF) social (NF) spatial synth speech AVG Adv (UA) Adv (US) AVG
Lang Model (WER) (NWERD) (WERD)
EN prkt-rnnt-1.1b 5.9 11.6 8.7 4.2 1.6 6.4 48.3 39.8 11.8 3.0 15.0 10.9 69.1 40.0
cnry-1b 6.0 13.9 18.2 4.4 2.7 15.0 45.7 36.7 15.3 5.7 17.5 14.8 61.7 38.3
prkt-ctc-1.1b 6.0 16.9 10.3 4.2 3.2 9.6 44.6 35.0 13.5 5.9 15.9 8.4 71.2 39.8
w2v2-lg-slf 7.7 41.0 30.7 13.1 17.6 20.3 69.5 67.6 26.3 15.2 33.5 7.0 41.0 24.0
wsp-md 7.9 12.4 27.8 2.8 3.3 3.5 41.8 35.5 5.2 6.1 15.4 3.7 56.6 30.1
wsp-lg 8.0 11.2 12.7 3.1 2.9 2.8 40.9 34.9 4.5 4.4 13.0 6.2 53.6 29.9
hubt-xl 8.4 38.9 29.1 15.5 16.2 20.8 69.5 71.2 26.4 13.5 33.5 13.9 36.8 25.3
wsp-bs 9.6 30.1 88.8 8.7 9.6 22.5 63.9 47.8 17.9 12.6 33.6 2.7 88.5 45.6
w2v2-lg 9.7 60.6 39.5 19.0 26.0 24.1 77.3 79.9 37.3 17.7 42.4 16.6 31.1 23.8
ES cnry-1b 3.2 699.9 21.7 17.4 7.6 30.5 - - 36.3 28.6 120.3 26.1 84.3 55.2
wsp-lg 5.8 6.8 19.4 12.3 5.5 9.2 - - 5.5 24.6 11.9 13.7 65.0 39.4
w2v2-lg-es 6.8 31.6 31.8 30.1 20.5 40.7 - - 89.0 104.1 49.7 33.9 71.0 52.4
wsp-bs 14.8 62.0 133.2 43.1 25.8 60.4 - - 58.1 87.8 67.2 19.5 159.5 89.5
$m$s-1b 15.7 26.3 27.9 32.7 9.3 43.3 - - 47.3 49.9 33.8 7.4 53.8 30.6
w2v2-bs-es 25.7 45.6 55.9 44.9 25.0 58.4 - - 103.3 152.7 69.4 10.0 33.8 21.9
Table 1: The utility and robustness of selected English and Spanish models (see Table LABEL:tab:full-results for more results). Utility is measured by WER of the models on clean speech. Robustness is measured by the NWERD on non-adversarially perturbed speech and WERD on adversarially perturbed speech. Adv (UA) refers to utterance agnostic attacks, while Adv (US) refers to utterance specific ones. The metrics are averaged over all datasets, perturbations, and severities in each category.
4.2 Robustness of ASR Models
Table 1 presents the utility and robustness of a subset of English and Spanish ASR models under non-adversarial and adversarial scenarios. The subset was selected to exclude small and/or less accurate models. The results of the excluded models, however, are used in § 4.3.
4.2.1 Robustness in Non-Adversarial Scenarios
English Models: In terms of average NWERD, we observe that wsp-lg emerges as the most robust model for non-adversarial scenarios, followed by prkt-rnnt-1.1b and wsp-md. Interestingly, cnry-1b, which is the top model on the Open ASR Leaderboard (OAL, Srivastav et al. 2023), ranks 5th on SRB. This result highlights the fact that SRB provides a more rigorous assessment of a model’s robustness than existing benchmarks like OAL. We also see that SRB reveals subtle weaknesses and strengths of various models.
For instance, we see that wsp-lg and wsp-md are significantly more robust to special effects (sFX) and spatial acoustics than other models, including cnry-1b.
@@ 4.2.1 Robustness in Non-Adversarial Scenarios
Figure 3: WERD of cnry-1b, wsp-md and wsp-lg on perturbations in the spatial acoustics and special effects categories.
@@
To identify the specific types of sFX and spatial acoustic perturbations against which cnry-1b lacks robustness, we plot the WERD on each perturbation within these categories (Fig. 3) we find that cnry-1b is much more sensitive than its peers to echo, real room impulse responses, and speed and pitch modifications. This analysis demonstrates that SRB can evaluate the robustness of ASR models at multiple granularities and can pinpoint the weaknesses of a given model. Detailed results can be found in Figs. 9 and 11 in the appendix. Given that no data augmentation, other than SpecAugment (Park et al., 2019), was used to train Whisper (Radford et al., 2023), this indicates that Whisper was trained on data that may have included digital media like music or movie soundtracks, and speech recorded in diverse acoustic environments – settings that may not be sufficiently represented in public data sources. Curating such diverse datasets is a promising direction for future work.
We also note that despite being pre-trained on 60K hours of speech, Wav2Vec and Hubert models severely lack robustness. Particularly concerning is their weak performance on accented speech, social scenarios and spatial acoustics, which models are very likely to encounter in the real world.
@@
Takeaways: (1) Despite topping the Open ASR Leaderboard, cnry-1b is significantly less robust than wsp-lg, which is ranked 10 on OAL. cnry-1b particularly lacks robustness to special effects and spatial acoustics. (2) Wav2Vec variants struggle against accented speech and social settings, thus, may not be suitable when users have diverse accents.
Spanish Models:We observe that wsp-lg is the most robust model against non-adversarial perturbations by some margin. We notice that all models, except wsp-lg, struggle against accented speech and yield high NWERS. cnry-1b is particularly, weak against accented speech with an NWERD of 700% (WERD=205%). Apart from accented speech, cnry-1b is quite robust on all other categories of non-adversarial perturbations. mms-1b is also fairly robust and, unlike other models, its NWERD does not vary erratically from one category to another.
4.2.2 Robustness in Adversarial Scenarios
@@
English models: We observe that w2v2-lg achieves the lowest WERD and thus is the most robust model against utterance-specific adversarial attacks. Interestingly, while Wav2Vec models exhibited mediocre robustness to non-adversarial perturbations, they are more robust to utterance-specific attacks, than Whisper, Canary, and Parakeet, which were the most robust on non-adversarial perturbations. We also note from Fig. 8 (in Appendix) that most Wav2Vec models are considerably more robust to attacks against TEDLIUM than against LibriSpeech, and the opposite is true for whisper and Canary models. Under utterance-agnostic attacks, the most robust models are mms-1b, wsp-bs, and wsp-sm. It is interesting to note that the smaller variants of Whisper limit the generalizability across utterances of the adversarial perturbations to a greater extent than their larger counterparts.
@@
Takeaway: Wav2Vec models are most robust to adversarial attacks; Models that are most robust to non-adversarial perturbations, are mediocre against adversarial perturbations; Canary and Parakeet models are highly vulnerable to utterance specific attacks.
@@
Spanish models: On Spanish, w2v2-bs-es is the most adversarially robust model. Generally, we observe that Wav2Vec models exhibit better robustness than Whisper and Canary under both utterance-agnostic and utterance-specific perturbations. This is similar to the trends observed in English speech (Fig. 8(c)). Detailed results can be found in Fig. 10 in the appendix.
Takeaway: General trends similar to English but WERD is higher when Spanish speech is attacked.
4.3 Correlates of Robustness
To glean insights that can inform future work, we have conducted the following analysis to model attributes that yield robust models. Specifically, we examine the impact of model size, architecture and accuracy, as well as training dataset size on robustness.
To determine if the prevailing practice of training DNNs with more parameters on larger datasets is yielding improvements in robustness, we use robust linear regression to fit a line to WERD vs. number of model parameters/size of the training data for the candidate models in Figs. 5(a) and 5(b), respectively. Increasing model size is correlated with improved robustness (lower WERD).
Figure 4: NWERD lineplot with non-adversarial and adversarial perturbations, three families of models.
To further isolate the impact of the model size we plot the NWERD of models from the same family in Fig. 4, which have similar architectures and training datasets. We note that larger models are more robust in the Whisper, Parakeet and Wav2Vec-2.0 families, but, surprisingly, not in the HuBert family.
Next, we consider the model architectures. The architectures of the models used in this paper can be divided in to three categories: sequence-to-sequence (seq2seq) models like Whisper and Canary, encoder only models trained with CTC loss (Graves et al., 2006) like the Wav2Vec family, and RNN-T models which are capable of streaming such as some variants of Parakeet. From Fig. 5(e) we see that in terms of non adversarial robustness RNN-T models outperform seq2seq and CTC models, but in terms of adversarial robustness CTC models achieve the lowest WERD.
We also measure the robustness-utility trade-off by plotting WERD and NWERD for adversarial and non-adversarial perturbations, respectively, against WER on clean data in Figs. 5(c) and 5(d). We observe that in both cases the relationship is positive, i.e. more accurate models tend to be more robust, however, the relationship between WERD on adversarial perturbations and clean WER is much weaker.
Finally, we measure the impact of training data size and, find that increasing training data appears to have only a minor influence on robustness (Fig. 5(b)).
Takeaway: (1) Larger models tend to be more robust, while smaller models, even if they are trained on large datasets, are less robust. This runs somewhat counter to the prevailing wisdom (Radford et al., 2023; Likhomanenko et al., 2020). (2) CTC models are more robust than seq2seq models to adversarial attack, but less robust than seq2seq and RNN-T models on non-adversarial perturbations. (3) Utility and robustness are positively correlated, but correlation is weaker for adversarial robustness.
Figure 5: (a & b) WERD for all models with robust regression fitted line on non-adversarial (blue) and adversarial (orange) perturbations, plotted against (a) number of parameters, and (b) hours of training data. (c % d) WERD and NWER on adversarial and non-adversarial perturbations are plotted against WER to illustrate the robustness-utility trade off. Pareto optimal points are highlighted. (e) Boxplot of WERD for models having various architectures.
4.4 Disparity in Robustness Across Population Sub-Groups
In the preceding analysis, we considered robustness aggregated over the entire population (i.e., dataset). However, populations are generally not homogeneous, and, thus, the robustness of the model may differ on various population sub-groups. Prior works have commonly analyzed sub-group fairness of ASR models by comparing the overall WER for each sub-group on a benchmark dataset (Koenecke et al., 2020). It is possible that models that are fair on average, may not be fair under certain conditions. In the following, we use SRB to uncover and analyze the disparities in the models’ robustness across four sub-groups: English and Spanish speech, and male and female speakers. We find that disparities indeed exist, with multi-lingual models generally being more robust for English than Spanish (Fig. 6), and most models being less robust for females than males.
4.4.1 Disparity in Robustness Across Languages in Multi-Lingual Models
Figure 6: Comparing the robustness of multi-lingual on English (solid) and Spanish (hatched).
@@ 4.4.1 Disparity in Robustness Across Languages in Multi-Lingual Models
We compare the robustness exhibited by multi-lingual models, wsp-lg, wsp-bs, cnry-1b and mms-1b on English and Spanish. The WERD of these models on both languages is presented in Fig. 6. We observe that Whisper models achieve lower WERD on English speech than on Spanish on almost all perturbation categories, while cnry-1b and mms-1b achieve similar WERD on some categories. We also note that the difference in WERD on some common perturbation categories, like environmental noise, and spatial acoustics, is much greater for wsp-lg than for cnry-1b.
@@
4.4.2 Disparity in Robustness Across Genders
To measure the disparity in transcription quality across genders (males/females), we compute the log of the ratio of the WERs of the ASR model on female and male speakers. We call this measure the Log WER Ratio (LWERR). A positive value of LWERR indicates that the model is biased against females and a negative value indicates that the model is biased against males.